Q. Which of the following function is periodic
a. sin x+sin √2x
b. sin 2x+cos 3x
c. e^xsin x
d. xsin x+cos x
1
Expert's answer
2019-07-09T12:32:53-0400

A function f: R → R is said to be periodic if there exists a number T > 0 such that f(x+T) = f(x) for all x ∈ R.

If there is a periodic function"f(x)"   with period T, the function "g(x) = f(kx)"   has of period:

"T^{\prime}=\frac T k"

a. sin x+sin √2x  is not periodic.


"y= \sin (x)" is a periodic function with period "2\pi".

"y= \sin (\sqrt{2}x)" is a periodic function with period "2\pi/\sqrt{2}=\sqrt{2}\pi"  

There is no integer, whose division by "2\pi" and "\sqrt{2}\pi" will give an integer.

Therefore, the function y = sin x+sin √2x is not periodic.

--------------------------------

b. sin 2x+cos 3x  is a periodic function with period "6\pi"

"y= \sin (2x)" is a periodic function with period "\pi" "(\frac {2\pi} {2}= \pi)" .

"y= \cos (3x)" is a periodic function with period "(\frac {2\pi} {3}= 2/3 \pi)" .


"y= \sin (2x)+\cos (3x)" is a periodic function with period "6\pi."

(the period is equal to the smallest number, with division of which by "\pi" and "2/3 \pi" we get integer numbers).

------------------------------------------------------

c. e^x sin x  is not periodic.

y=sin(x) is a periodic function with period "2\pi."

y= e^x is not periodic.

The product of functions e^x and sin x will not be a periodic function.


-----------------------------------------

d. xsin x+cos x is not periodic.


y=xsin x is not periodic. (y=x -  is not periodic, y=sin x -  is periodic "2\pi." )

y=cos x is a  periodic function with period "2\pi." . .

y= xsin x+cos x  is not periodic. (The sum of the non-periodic (y=xsin x ) and periodic (y=cos x) functions is not a periodic function).